What quantity of 60% acid solution must be mixed with 30% solution to produce 300mL of a 50% solution?
Answer by Nate(3500)

(x(60%) + (300 - x)(30%))/(300) = 50%
(x(0.6) + (300 - x)(0.3))/(300) = 0.5
0.6x + 90 - 0.3x = 150
0.3x = 60
x = 200
200mL of 60% and 100mL of 30%